Circular No. 5638

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992bd IN NGC 1097 Chris Smith, Cerro Tololo Interamerican Observatory; and Lisa Wells,

Kitt Peak National Observatory, report their discovery of a supernova located 1′′.5 east and 9′′ south of the nucleus of NGC 1097 (α = 2^{h}44^{m}11^{s}, δ = −30^{o}29′.1, equinox 1950.0). The object was found in a 20-min I-band CCD exposure taken by Smith on Oct. 12.26 UT with the CTIO 0.9-m telescope; it is also apparent in B, V, and R CCD images taken subsequently, and has a magnitude of V ∼ 15 and a color of B–V ∼ +2.6. Nothing is visible at this location in published images of the nucleus of NGC 1097 (e.g., Rickard 1975, A.Ap. 40, 339). Due to the highly reddened nature of this object, infrared and radio observations may prove interesting.

M. Della Valle and P. Bouchet, European Southern Observatory, report: “Inspection of a spectrogram (range 400-830 nm, resolution ∼ 1 nm) of SN 1992bd, obtained under poor meteorological conditions on Oct. 17.1 UT with the New Technology Telescope (+ EMMI) at La Silla, shows this object to be a type-II supernova. The spectrum is dominated by the emissions of Hα and Na I D, flanked by deep P-Cyg profiles. The expansion velocity, corresponding to the difference between the rest wavelength in the galaxy’s frame and the minimum of the absorption, is ∼ 7500 km/s. Other strong emission lines are observed at 489 (Hβ), 493, 543, 570, and 615 nm.”

NOVA CYGNI 1992 G. S. Rossano, R. J. Rudy, R. C. Puetter, and D. K. Lynch, Space and

Environment Technology Center, The Aerospace Corporation, report: “We observed this object on Oct. 6.3 and 8.3 UT with the Lick Observatory 3-m Shane telescope (+ cooled scanning germanium spectrometer with a 10′′ × 8′′ oval aperture). Spectroscopic observations (range 0.75-1.35 μm, resolution 300-550) indicate that N Cyg 1992 has changed little in appearance since August, showing strong lines of the Paschen series, and strong He I 1083.0-nm, along with other lines of He I, He II, O I, [O II], [N I], [S II], and [S III]. The strength of the O I lines has decreased, and an unidentified feature at 1110 nm has increased markedly in strength. A monochromatic magnitude of 8.9 at 1.25 μm was observed on Aug. 8.3.”

Selected visual magnitude estimates: Sept. 23.24 UT, 8.8 (R. A. Keen, Golden, CO); 30.14, 9.4 (W. G. Dillon, Missouri City, TX); Oct. 4.79, 9.1 (H. Dahle, Oslo, Norway); 9.01, 9.1 (B. H. Granslo, Fjellhamar, Norway); 13.94, 9.4 (P. Schmeer, Bischmisheim, Germany).

1992 October 17 Daniel W. E. Green